OpenStudy (anonymous):

In 1990 the population of the U.S. was approximately 250 million and was expected to grow according to the function P(x) = 250,000,000e 0.009x, where x is the number of years after 1990. When will there be three times as many people in the U.S. as in 1990, according to this model?

OpenStudy (stormfire1):

\[750,000,000=250,000,000*e^{0.009x}\] \[3 = e^{.009x}\] \[\ln 3 = 0.009x\] x = ~122 years